Kendra has $1.75 in dimes and nickels. If she has 3 times as many dimes as nickels, how many of each coin does she have?
Answer by addingup(3677)

d = 3n
0.10d+0.05n = 1.75
0.10(3n)+0.05n = 1.75
0.30n+0.05n = 1.75
0.35n = 1.75
n = 5 and she has 3 times as many dimes:
d = 5*3 = 15
-----------------------------
check:
5*0.05+15*0.10 = 0.25+1.50 = 1.75 Correct.
J